### v1

# A Lightweight Framework for GPU-Accelerated Robot Learning github.com/mujocolab/mjlab

Abstract
We present mjlab, a lightweight, open-source framework for robot learning that combines GPU-accelerated simulation with composable environments and minimal setup friction. mjlab adopts the manager-based API introduced by Isaac Lab (Mittal et al., 2025), where users compose modular building blocks for observations, rewards, and events, and pairs it with MuJoCo Warp (Google DeepMind and NVIDIA, 2025) for GPU-accelerated physics. The result is a framework installable with a single command, requiring minimal dependencies, and providing direct access to native MuJoCo data structures. mjlab ships with reference implementations of velocity tracking, motion imitation, and manipulation tasks.
Figure 1: Unitree G1 humanoids performing a dance trained with the motion tracking pipeline from BeyondMimic (Liao et al., 2025). Thousands of parallel environments are simulated on a single GPU and visualized in mjlab’s web-based Viser viewer.
1  Introduction
Reinforcement learning has become a powerful tool for training robot controllers in simulation and transferring them to real hardware. The fidelity of this sim-to-real pipeline hinges on getting simulation details right: actuator dynamics, contact modeling, sensor noise, and domain randomization all require careful implementation. Middleware that handles this infrastructure, ideally written and vetted by the people closest to the simulator, frees researchers to focus on what matters most: shaping reward functions, designing curriculums, and iterating on policies.
Several frameworks address this need. Isaac Lab (Mittal et al., 2025) provides a comprehensive, GPU-accelerated platform with a rich manager-based API for composing RL environments. However, it requires the Omniverse runtime, which adds installation complexity and startup latency. Its physics engine, PhysX, was closed-source until recently, making low-level debugging and introspection difficult. MuJoCo Playground (Zakka et al., 2025) takes the opposite approach: minimal abstractions and monolithic environment definitions that are easy to hack and quick to prototype. This design excels for one-off experiments, but code duplication across robots and tasks accumulates quickly, making multi-robot or multi-task codebases difficult to maintain. There remains a gap for a framework that is both lightweight and built on a proven orchestration API with access to best-in-class physics.
mjlab fills this gap. It adopts Isaac Lab’s manager-based design, where users compose self-contained building blocks for observations, rewards, events, and commands, and pairs it with MuJoCo Warp for GPU-accelerated physics simulation. The result is a framework with minimal dependencies, fast startup, direct access to native MuJoCo model and data structures, and a PyTorch (Ansel et al., 2024)-native interface for policy training. mjlab ships with three robot morphologies and three reference tasks. The remainder of this report describes the design philosophy (Section 2), architecture (Section 3), core components (Section 4), the manager-based API (Section 5), shipped tasks (Section 6), and software design (Section 7).
2  Design Philosophy and Scope
mjlab is designed around three core engineering commitments: (1) minimal installation friction, (2) transparent and inspectable physics, and (3) tight integration with the MuJoCo ecosystem. These commitments motivate deliberate trade-offs in implementation and scope, which we make explicit below.
Physics backend.
mjlab targets a single physics stack, MuJoCo Warp, to prioritize simulation transparency and debuggability. The framework exposes MuJoCo-native MjModel and MjData structures for direct inspection and state access. A non-goal is cross-simulator portability (Amazon FAR et al., ); mjlab favors precise control and interpretability over backend generality.
Sensing and rendering.
mjlab provides depth and raycast sensors for geometric perception. High-fidelity RGB rendering is out of scope. This does not preclude vision-based policies: a common approach is to train privileged policies using full state, then distill into vision-based controllers using external rendering.
Environment composition.
mjlab adopts Isaac Lab’s manager-based API, where environments are built from reusable, composable components rather than monolithic definitions. Modular terms for observations, rewards, events, and curricula reduce code duplication across tasks. The implementation remains fully MuJoCo-native; users work directly with MuJoCo models and conventions.
Scope and extensibility.
mjlab provides infrastructure for rigid-body robot learning that is intended to be extended to custom robots, tasks, sensors, and actuators. The framework includes three example robots and tasks that demonstrate use of the API.
Together, these choices reflect a focus on fast iteration, transparent physics, and MuJoCo-native workflows, rather than feature completeness or broad simulator support.
3  Architecture
Entity 1(G1 robot)MJCF + configsEntity 2(Cube)MJCFEntity 3(Terrain)PythonMjSpecMjModelMuJoCo Warp(GPU, NN worlds)ManagerBasedRlEnvObsActionRewardTermEventCmdCurricRSL-RLTrained Policy
Figure 2: Entities are composed into an MjSpec, compiled, and transferred to MuJoCo Warp for GPU simulation. The ManagerBasedRlEnv orchestrates the MDP; RSL-RL handles training.
MuJoCo Warp. MuJoCo Warp Google DeepMind and NVIDIA (2025) (docs) is a GPU-accelerated backend for MuJoCo built on NVIDIA Warp (Macklin, 2022). It preserves MuJoCo’s (Todorov et al., 2012) familiar MjModel/MjData paradigm: MjModel holds the static kinematic and dynamic description of the scene, while MjData carries the time-varying simulation state. The key addition is a leading world dimension: a single MjData object holds the state of $N$ independent simulation instances in parallel, enabling thousands of environments to be stepped in parallel. Model parameters are shared across all worlds by default, and individual fields can be expanded to vary per-world when domain randomization requires it. mjlab further captures the simulation step as a CUDA graph: the kernel execution sequence is recorded once and replayed on subsequent calls, eliminating CPU-side dispatch overhead.
Scene pipeline. mjlab constructs scenes by composing entity descriptions defined via MJCF into a single MjSpec. This specification is compiled into an MjModel on the CPU, then transferred to the GPU via MuJoCo Warp’s put_model and put_data routines.
Manager-based paradigm. On top of the simulation layer, mjlab adopts the manager-based environment design introduced by Isaac Lab. Users define their environment by composing small, self-contained terms—reward functions, observation computations, domain randomization events—and register them with the appropriate manager. Each manager handles the lifecycle of its terms: calling them at the right point in the simulation loop, aggregating their outputs, and exposing diagnostics. Terms can be plain functions for stateless computations or classes that inherit from ManagerTermBase when they need to cache expensive setup (e.g., resolving regex patterns to joint indices at initialization) or maintain state across steps with a per-episode reset() hook. This abstraction introduces overhead compared to writing monolithic step functions, but benefits in modularity, testability, and rapid iteration outweigh the cost—a lesson reinforced by our experience with the more minimal MuJoCo Playground API, where the absence of such structure made environments harder to maintain and extend.
4  Components
mjlab provides a set of core components that managers operate on during the simulation lifecycle.
Entities. An entity represents a physical object in the scene, such as robots, manipulated objects, or static fixtures. While Isaac Lab splits this concept across multiple specialized classes (Articulation, RigidObject, DeformableObject, RigidObjectCollection, and others), mjlab uses a single Entity class that covers all variations through two independent properties: base type (fixed or floating) and articulation (with or without joints). Runtime queries such as is_fixed_base and is_articulated let downstream code branch where needed without requiring a class hierarchy. Each entity maintains an EntityData structure that aggregates kinematic quantities such as body positions, orientations, and velocities, making them available to observation and reward terms without redundant computation.
Sensors. Sensors in mjlab follow a layered design. Users can wrap sensors already defined in their MJCF files and read them through MuJoCo’s native sensordata buffer, or define them entirely in Python configuration. On top of this native layer, mjlab provides custom sensor types that go beyond what MuJoCo offers: a ray-cast sensor for terrain height scanning, a contact sensor that tracks collision forces between body pairs with air-time and force-history bookkeeping, and an experimental tiled-rendering camera. All sensors implement a common interface with per-step caching and optional debug visualization.
Actuators. Actuators follow the same layered philosophy. Users can wrap actuators defined in MJCF—preserving the exact behavior specified in the model file—or create MuJoCo-native actuator types (motor, position, velocity, muscle) from configuration. Beyond these native options, mjlab provides custom implementations that compute torques on the GPU outside of MuJoCo: an ideal PD controller, a DC motor model with velocity-dependent torque saturation, and a learned MLP actuator for capturing hardware-specific dynamics from data. All three tiers coexist within the same entity. Actuation delay—common in real robots—is modeled by a wrapper class that buffers control signals and replays them with a latency quantized to the physics timestep.
Terrain. The terrain module generates ground geometry for locomotion training. Users can spawn a flat ground plane or compose a grid of sub-terrain patches, each drawn from a library of terrain types. Two families are provided: primitive terrains built from MuJoCo box geoms (flat surfaces, pyramid stairs, random-height grids) and heightfield terrains for smoother, continuous profiles (sloped pyramids, uniform noise, sinusoidal waves). Every terrain type accepts a difficulty parameter that scales its features—step height, slope angle, noise amplitude—from trivial to challenging. The grid generator supports two modes: random, where each patch is sampled independently, and curriculum, where rows encode increasing difficulty so that the curriculum manager can advance robots to harder terrain as performance improves. Users can subclass the base terrain configuration to define custom terrain types (Figure 3).
Figure 3: A composite terrain grid generated by mjlab. Sub-terrain patches include flat ground, stairs, sloped pyramids, heightfield noise, and wave patterns. Difficulty increases along one axis for curriculum-based training.
Viewer. mjlab supports two visualization backends. The native MuJoCo viewer provides the full simulate experience—pause and resume, contact point and force visualization, interactive perturbations, and more. For headless servers or remote machines, mjlab offers a web-based viewer built on Viser (Yi et al., 2025) that supports pause, resume, and contact visualization in the browser.
5  Manager-Based API
mjlab environments implement the Gym (Brockman et al., 2016) interface, a standard API for defining Markov decision processes (MDPs). Environments expose two methods: reset() initializes a new episode and returns the first observation, while step() applies an action and returns the next observation, reward, and termination signal. Each call to step() proceeds through a fixed pipeline:
  1. 1.
    Act. The action manager updates the action history and splits the policy output into per-term slices.
  2. 2.
    Simulate. For each of $d$ decimation sub-steps: apply actuator commands, write controls to the simulation, advance physics, and update entity state.
  3. 3.
    Terminate. The termination manager evaluates stop conditions and flags environments for reset.
  4. 4.
    Reward. The reward manager computes and aggregates weighted reward terms.
  5. 5.
    Reset. Flagged environments are reset: the curriculum manager updates difficulty, entity states are reinitialized, and the event manager applies reset hooks.
  6. 6.
    Command. The command manager generates or resamples goal signals.
  7. 7.
    Apply Events. The event manager applies periodic events (e.g., external pushes).
  8. 8.
    Observe. The observation manager computes observations for the next policy query.
This manager-based orchestration follows the design introduced by Isaac Lab (Mittal et al., 2025), where each manager encapsulates a specific aspect of the MDP and users compose environments by registering modular terms.
Action manager. The action manager receives the policy’s output tensor, splits it according to the registered action terms, and routes each slice to the appropriate entity’s actuators. It supports action clipping and tracks the two most recent actions, which observation and reward terms can reference (e.g., to compute an action rate penalty).
Termination manager. The termination manager evaluates a set of Boolean conditions each step—contact with illegal bodies, joint limit violations, timeout—and flags environments for reset. It also monitors simulation state for NaN and Inf values, catching numerical instabilities before they propagate. When an instability is detected, mjlab captures a rolling buffer of recent states that users can replay in an interactive viewer to diagnose the failure. The observation and reward managers additionally report which specific terms produced invalid values, narrowing the search from “something exploded” to the exact offending computation.
Reward manager. Reward terms are functions or stateful classes that return per-environment scalars. The manager computes a weighted sum of all active terms, automatically scaling each by the environment’s control timestep to make reward magnitudes invariant to the simulation frequency. Per-term episodic sums are tracked and logged, making it straightforward to diagnose which terms dominate training. Terms that produce NaN or Inf values are detected and reported individually.
Curriculum manager. The curriculum manager adjusts training conditions based on policy performance. Examples include scaling reward term weights, widening command velocity ranges, or advancing robots to harder terrain rows as success rates improve.
Event manager. Events are hooks that execute at specific points in the environment lifecycle: at startup, on reset, or at fixed intervals during training. The most common use case is domain randomization. When an event modifies a field of the MuJoCo model (e.g., friction coefficients or link masses), mjlab expands that field from a shared value to a per-world array and rebuilds the CUDA capture graph to reflect the new memory layout. This happens transparently to the user, who simply specifies properties and ranges to randomize over.
Command manager. The command manager generates goal signals for the policy—velocity targets, tracking waypoints, or object poses—and resamples them at configurable intervals. Commands are exposed to the observation manager so the policy can condition on the current goal.
Observation manager. Observation terms are composable functions that read from entity data, sensors, or commands and return tensors. The manager chains these terms through a configurable pipeline: raw values are optionally clipped, scaled, corrupted with noise, delayed by a configurable number of control steps, and concatenated with a history buffer. Multiple observation groups (e.g., policy and critic) can coexist, each with its own processing pipeline, enabling asymmetric actor-critic architectures.
6  Tasks
Figure 4: The three robot morphologies shipped with mjlab: Unitree G1 humanoid (left), Unitree Go1 quadruped (center), and YAM robot arm (right). All models are adapted from MuJoCo Menagerie (Zakka et al., 2022).
mjlab ships with three reference tasks spanning locomotion, whole-body control, and manipulation. The codebase is intentionally lean: three robot morphologies—a humanoid (Unitree G1), a quadruped (Unitree Go1), and a robot arm (YAM)—cover a broad range of research use cases while keeping the repository easy to navigate and maintain.
6.1  Locomotion: Velocity Tracking
The velocity tracking task trains a locomotion controller to follow linear and angular velocity commands on flat or rough terrain. The policy observes IMU readings, projected gravity, joint positions and velocities, the previous action, and the commanded twist. Reward terms encourage accurate velocity tracking while penalizing excessive body angular velocity, angular momentum, joint limit violations, action rate, and foot slip. Terrain variants range from flat ground to composite grids of stairs and rough heightfields, with terrain difficulty tied to the curriculum manager. Videos show a natural running gait on the G1 humanoid emerging from this reward structure in both simulation [link] and on real hardware [link].
6.2  Whole-Body Control: Motion Imitation
The motion imitation task trains a humanoid (Unitree G1) to track reference motion clips, implementing the DeepMimic (Peng et al., 2018) framework with extensions from BeyondMimic (Liao et al., 2025). The policy observes an anchor pose from the reference trajectory, base velocities, joint states, and the current action. Reward terms penalize deviations in global root position and orientation, relative body poses, and body velocities. A self-collision cost discourages interpenetration. A demonstration of a trained triple spinkick is available online [link].
6.3  Manipulation: Cube Lifting
The manipulation task trains a YAM robot arm to lift a cube to a target pose. The policy observes joint positions and velocities, the vector from end-effector to cube, the cube-to-goal error, and the previous action. A staged reward structure first guides the end-effector toward the cube, then rewards lifting the cube to the target height. Contact sensors on the end-effector and ground plane provide auxiliary signals. A video demonstration is available online [link].
7  Software Design
A framework’s long-term usability depends as much on how its code is organized as on what it can simulate. We made several deliberate departures from Isaac Lab’s software patterns to reduce boilerplate, catch errors earlier, and lower the barrier to writing new tasks.
PyTorch-native interface. MuJoCo Warp stores simulation state in Warp arrays, while policy training typically happens in PyTorch. mjlab bridges this gap with a TorchArray abstraction: a zero-copy wrapper that exposes Warp arrays as PyTorch tensors. Users write rewards, observations, and custom logic in pure PyTorch without touching Warp kernels, and the framework handles memory sharing transparently.
Instance-based configuration. Isaac Lab organizes environment parameters through deeply nested dataclass hierarchies: each new task inherits from a base config class and overrides fields via __post_init__ mutations or field redeclarations. This pattern is fragile: __post_init__ overrides can silently ignore constructor arguments, and a misspelled field name creates a new field rather than overriding the parent’s, with no error at any stage. mjlab replaces class-based configs with plain config instances. Observation, reward, event, and other term configurations are held in typed dictionaries (dict[str, RewardTermCfg], etc.) rather than in bespoke dataclass subclasses. Creating a task variant is a matter of copying an existing config and mutating it, not defining a new class. This eliminates the inheritance pitfalls while making configs easier to read, diff, and compose programmatically.
CLI-first configuration. Because configs are ordinary dataclass instances with typed fields, mjlab can expose every parameter at the command line automatically via tyro. Users override settings like reward weights, observation noise, simulation timestep with a single flag, without writing configuration files or subclassing.
Co-located definitions. Isaac Lab separates configuration dataclasses from their implementations: all manager term configs live in a single manager_term_cfg.py, while the corresponding classes are spread across other files. Configs must reference their implementation via a class_type field, adding indirection. mjlab keeps each config dataclass in the same file as its implementation—ActionTermCfg alongside ActionTerm, for example—so related code is always collocated and navigation is straightforward.
Static typing and testing. The entire codebase is statically type-checked with both pyright and ty, and ships a py.typed marker for downstream consumers. A test suite of over 40 files covers managers, sensors, actuators, domain randomization, task execution, and more.
AI-assisted development. These design choices—typed configs, comprehensive tests, and a lean codebase—make mjlab particularly amenable to AI coding agents. Tools such as Claude Code can type-check their own edits, run the test suite to verify correctness, and iterate autonomously with minimal human intervention. In practice, several recent contributions were implemented end-to-end by AI agents (e.g., PR #532), suggesting that a well-structured codebase lowers the barrier for not only human contributions but also for automated ones.
Minimal dependencies. A guiding constraint throughout mjlab’s development was keeping installation friction-free. The framework avoids heavyweight dependencies, and is designed around uv (Astral, 2025), a fast Python package manager that eliminates explicit environment setup. Users can go from a fresh clone to a running training job in seconds:
git clone https://github.com/mujocolab/mjlab && cd mjlab
uv run train Mjlab-Velocity-Flat-Unitree-G1 \
  --env.scene.num-envs 4096 \
  --agent.max-iterations 10_000
Environment creation, installation, and dependency resolution happen automatically. Training uses RSL-RL (Schwarke et al., 2025) for on-policy algorithms, with multi-GPU scaling available via torchrunx (Khandelwal, Apoorv and Curtin, Peter, 2025).
8  Adoption
mjlab has been deployed in UC Berkeley’s graduate robotics course (ME 292b/193b), where students new to reinforcement learning trained their first legged locomotion policies. The framework has been adopted by various open-source projects [1, 2] and featured in a tutorial from a popular YouTuber with over 1.5M subscribers [link].
9  Acknowledgments
We would like to thank Yuval Tassa, Erik Frey, Baruch Tabanpour, and Taylor Howell for their help with MuJoCo and MuJoCo Warp. We also thank Adam Allevato, Grégoire Passault, Anoop Gadhrri, Sergi de la Muelas Acosta, Mustafa Haiderbhai, Haoyang Wang, Alphonsus Adu-Bredu, Bolaji Adesina, Tatsuki Tsujimoto, Vittorio Caggiano, Zahi Kakish, and Andrew Carr for their code contributions, and Robin Deits for finding and reporting bugs.

### v2

@@
 Sensing and rendering.
-mjlab provides depth and raycast sensors for geometric perception.
-High-fidelity RGB rendering is out of scope.
+mjlab provides ray-cast, depth, and RGB camera sensors for geometric and visual perception.
+Photorealistic rendering is out of scope.
 This does not preclude vision-based policies: a common approach is to train privileged policies using full state, then distill into vision-based controllers using external rendering.
@@
 3  Architecture
-Entity 1(G1 robot)MJCF + configsEntity 2(Cube)MJCFEntity 3(Terrain)PythonMjSpecMjModelMuJoCo Warp(GPU, NN worlds)ManagerBasedRlEnvObsActionRewardTermEventCmdCurricRSL-RLTrained Policy
+Entity 1(G1 robot)MJCF + configsEntity 2(Cube)MJCFEntity 3(Terrain)PythonMjSpecMjModelMuJoCo Warp(GPU, NN worlds)ManagerBasedRlEnvObsActionRewardTermEventCmdCurricMetricsRSL-RLTrained Policy
 Figure 2: Entities are composed into an MjSpec, compiled, and transferred to MuJoCo Warp for GPU simulation. The ManagerBasedRlEnv orchestrates the MDP; RSL-RL handles training.
 MuJoCo Warp.
-MuJoCo Warp Google DeepMind and NVIDIA (2025) (docs) is a GPU-accelerated backend for MuJoCo built on NVIDIA Warp (Macklin, 2022).
+MuJoCo Warp (Google DeepMind and NVIDIA, 2025) (docs) is a GPU-accelerated backend for MuJoCo built on NVIDIA Warp (Macklin, 2022).
 It preserves MuJoCo’s (Todorov et al., 2012) familiar MjModel/MjData paradigm: MjModel holds the static kinematic and dynamic description of the scene, while MjData carries the time-varying simulation state.
@@
 Terms can be plain functions for stateless computations or classes that inherit from ManagerTermBase when they need to cache expensive setup (e.g., resolving regex patterns to joint indices at initialization) or maintain state across steps with a per-episode reset() hook.
-This abstraction introduces overhead compared to writing monolithic step functions, but benefits in modularity, testability, and rapid iteration outweigh the cost—a lesson reinforced by our experience with the more minimal MuJoCo Playground API, where the absence of such structure made environments harder to maintain and extend.
+This abstraction introduces overhead compared to writing monolithic step functions, but benefits in modularity, testability, and rapid iteration outweigh the cost, a lesson reinforced by our experience with the more minimal MuJoCo Playground API, where the absence of such structure made environments harder to maintain and extend.
 4  Components
@@
 Users can wrap sensors already defined in their MJCF files and read them through MuJoCo’s native sensordata buffer, or define them entirely in Python configuration.
-On top of this native layer, mjlab provides custom sensor types that go beyond what MuJoCo offers: a ray-cast sensor for terrain height scanning, a contact sensor that tracks collision forces between body pairs with air-time and force-history bookkeeping, and an experimental tiled-rendering camera.
+On top of this native layer, mjlab provides custom sensor types that go beyond what MuJoCo offers: a BVH-accelerated ray-cast sensor that supports grid and pinhole ray patterns and returns per-ray distances, hit positions, and surface normals; a contact sensor that tracks collision forces between body pairs with air-time and force-history bookkeeping; and a GPU-accelerated camera sensor that renders RGB and depth across all worlds in parallel and supports both perspective and orthographic projection.
 All sensors implement a common interface with per-step caching and optional debug visualization.
@@
 Actuators follow the same layered philosophy.
-Users can wrap actuators defined in MJCF—preserving the exact behavior specified in the model file—or create MuJoCo-native actuator types (motor, position, velocity, muscle) from configuration.
+Users can wrap actuators defined in MJCF, preserving the exact behavior specified in the model file, or create MuJoCo-native actuator types (motor, position, velocity, muscle) from configuration.
 Beyond these native options, mjlab provides custom implementations that compute torques on the GPU outside of MuJoCo: an ideal PD controller, a DC motor model with velocity-dependent torque saturation, and a learned MLP actuator for capturing hardware-specific dynamics from data.
 All three tiers coexist within the same entity.
-Actuation delay—common in real robots—is modeled by a wrapper class that buffers control signals and replays them with a latency quantized to the physics timestep.
+Actuation delay, common in real robots, is modeled by a wrapper class that buffers control signals and replays them with a latency quantized to the physics timestep.
 Terrain.
@@
 Users can spawn a flat ground plane or compose a grid of sub-terrain patches, each drawn from a library of terrain types.
-Two families are provided: primitive terrains built from MuJoCo box geoms (flat surfaces, pyramid stairs, random-height grids) and heightfield terrains for smoother, continuous profiles (sloped pyramids, uniform noise, sinusoidal waves).
+Two families are provided: primitive terrains built from MuJoCo box geoms (flat surfaces, stairs, stepping stones, narrow beams, and random-height grids) and heightfield terrains for smoother, continuous profiles (sloped pyramids, Perlin noise, and sinusoidal waves).
 Every terrain type accepts a difficulty parameter that scales its features—step height, slope angle, noise amplitude—from trivial to challenging.
@@
 It supports action clipping and tracks the two most recent actions, which observation and reward terms can reference (e.g., to compute an action rate penalty).
+Shipped action terms include joint-space position and velocity targets as well as a differential inverse kinematics term that converts task-space pose commands into joint displacements via damped least-squares, with optional null-space posture regularization and soft joint-limit avoidance.
 Termination manager.
@@
 The most common use case is domain randomization.
-When an event modifies a field of the MuJoCo model (e.g., friction coefficients or link masses), mjlab expands that field from a shared value to a per-world array and rebuilds the CUDA capture graph to reflect the new memory layout.
-This happens transparently to the user, who simply specifies properties and ranges to randomize over.
+mjlab provides a domain randomization module covering body, joint, geometry, camera, light, material, and tendon parameters, each with configurable operations (absolute, additive, multiplicative) and sampling distributions (uniform, log-uniform, Gaussian).
+When an event modifies a field of the MuJoCo model, mjlab expands that field from a shared value to a per-world array and rebuilds the CUDA capture graph to reflect the new memory layout.
+For inertial parameters, the module implements physically consistent randomization based on the pseudo-inertia parameterization of Rucker and Wensing (2022), guaranteeing that randomized mass, center of mass, and rotational inertia remain fully physically consistent for any perturbation magnitude.
+The breadth of randomizable parameters, from link geometry and inertial properties to joint dynamics and actuator gains, combined with physical consistency, makes the module applicable beyond sim-to-real transfer to problems such as system identification and hardware co-design.
 Command manager.
@@
 Multiple observation groups (e.g., policy and critic) can coexist, each with its own processing pipeline, enabling asymmetric actor-critic architectures.
+Metrics manager.
+The metrics manager computes and logs custom per-episode metrics that are not tied to any specific manager, such as task success rates or tracking errors, providing a unified interface for monitoring training progress.
 6  Tasks